OpenStudy (anonymous):

Three resistors of 10.0 W, 20.0 W, and 25.0 W are connected in parallel across a 100-V battery. What is the equivalent resistance of the circuit and the current flowing through the 10.0-W resistor?

OpenStudy (anonymous):

change the watt ratings to ohms for the resistors: 10w = 100V * amp I= amp = 0.1 P = I^2 * R 10w = (0.1)^2 *R R 1 = 1000 ohm 20w = 100V * I I = 0.2 amp 20w = (0.2)^2* R2 R2 = 500 ohm 25w = 100V *I I = (0.25)^2 * R3 R3 = 400 ohm For parallel resistors: 1 / R eqiv = 1/ R1 + 1/R2 + 1/ R3 1 / R eqiv = 1 / 1000 + 1/ 500 + 1/ 400 R equivalence = 2000 / 11 = 182 ohms The current flowing through the 10.0-W resistor? 10w = 1000 ohm V = IR 100V = I * 1000 ohm I = 0.1 amp current thru 10 watt resistor

OpenStudy (radar):

There are many ways to get these two things (equivalent resistance, and current through the 10 watt resistor. For the equivalent resistance. Total power is 10 + 20 + 25 watts or 55 watts P = (E^2)/R or 55=(100^2)/R Solving for R: R=(100^2)/55 = 181.81818 or 181.2 Ohms The second thing current thru the 10 W resistor. P=EI 10 = 100 I or I = 10/100 =.1 amps

OpenStudy (radar):

The solution by biire2u is correct and done in a very good step by step manner. I just wanted to show that it was unnecessary to solve for the individual resistances in order to solve this problem.